OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x2 + 28x - 8 A. Minimum Value: 7 Range y > 7 B. Minimum Value: -106 Range y > -7 C. Minimum Value: -7 Range y > -7 D. Minimum Value: -106 Range y > -106 not "B"

OpenStudy (anonymous):

f(x) = 2x2 + 28x - 8 =2(x^2+14x-4) =2(x^2+14x+49-49-4) =2(x+7)^2+2(-53) =2(x+7)^2-106 It is B

OpenStudy (anonymous):

I mean D

OpenStudy (anonymous):

Yes the answer is D.